# LOCAL: Learning with Orientation Matrix to Infer Causal Structure from Time Series Data

### v2

@@
Abstract.
Discovering the underlying Directed Acyclic Graph (DAG) from time series observational data is highly challenging due to the dynamic nature and complex nonlinear interactions between variables. Existing methods often struggle with inefficiency and the handling of high-dimensional data. To address these research gap, we propose LOCAL, a highly efficient, easy-to-implement, and constraint-free method for recovering dynamic causal structures. LOCAL is the first attempt to formulate a quasi-maximum likelihood-based score function for learning the dynamic DAG equivalent to the ground truth. On this basis, we propose two adaptive modules for enhancing the algebraic characterization of acyclicity with new capabilities: Asymptotic Causal Mask Learning (ACML) and Dynamic Graph Parameter Learning (DGPL). ACML generates causal masks using learnable priority vectors and the Gumbel-Sigmoid function, ensuring the creation of DAGs while optimizing computational efficiency. DGPL transforms causal learning into decomposed matrix products, capturing the dynamic causal structure of high-dimensional data and enhancing interpretability. Extensive experiments on synthetic and real-world datasets demonstrate that LOCAL significantly outperforms existing methods, and highlight LOCAL’s potential as a robust and efficient method for dynamic causal discovery. Our code will be available soon.
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Causal reasoning and diagnostics; Temporal reasoning
1. Introduction
@@
Exploration of the underlying causal generation process of dynamic systems is an important task (Cheng et al., 2024a; Gong et al., 2024) for trustworthy machine learning. Unfortunately, it is unethical, impossible due to technical reasons, or expensive to conduct intervention experiments on the dynamic systems of certain domains (Li et al., 2023a; Cai et al., 2023a). Another challenge is to infer about the structure which may be high dimensional and nonlinear. Some recent works (Pamfil et al., 2020; Sun et al., 2023; Gao et al., 2022; Fan et al., 2023) have made significant efforts by employing dynamic Bayesian networks (DBNs) with observational and interventional data: among dynamic systems, as illustrated in Figure 1. The variable $x_{i}$ at timestep $t$ is affected by which variables $x_{j}$ at the same time step (instantaneous dependency) and which variables $x_{j}$ at the previous timestep (lagged dependency)? This question highlights the crucial roles of those algorithms in the interpretable performance of the trained models.
@@
Figure 1. Illustration of instantaneous dependency (solid lines) and lagged dependency (dashed lines) dependencies in a DBN with $d=3$ nodes and autoregression order $p=2$. For clarity, we display edges that do not influence the variables at time $t$ in a lighter shade.
@@
In order to study the nonparametric DBN, DYNOTEARS (Pamfil et al., 2020) (i.e., a score-based approach to learning DBNs with a differentiable DAG constraint (Zheng et al., 2018)) was proposed as a proxy to capture the learn the parents of child variables. However, in Section 4.2, our practical analysis shows that the DYNOTEARS algorithm and its extensions (Sun et al., 2023; Gao et al., 2022; Fan et al., 2023) adopting matrix exponential constraints require an extremely long time to optimize in high-dimensional dynamic systems, even if they smartly adopt interventional data to enhance the identifiability (Li et al., 2023b, 2024). Then, it is natural to need a proxy model that can fasterly infer dynamic causal structures in high-dimensional situations, which is also the main goal of our work.
@@
Recently, much of the research on dynamic causal structure learning has concentrated on applying soft sparsity and DAG constraints. For instance, Golem (Ng et al., 2020) formulated a likelihood-based score function for handling the causality of thousands of nodes. Yu et al. (Yu et al., 2023) extended it for recovery dynamic causal structure. Concurrently, (Fang et al., 2024) verified the feasibility of further accelerating the learning of DAG based on this likelihood function both theoretically and experimentally. These approaches are aimed at enhancing flexibility and scalability in high-dimensional and circumventing rigid structural constraints.
@@
Building on these insights, we propose a novel framework for Learning with Orientation matrix to infer CAusaL structure from time series data, which we call LOCAL. We develop a quasi-maximum likelihood-based dynamic structure learning method with identifiability guarantee. Powered by this quasi-maximum likelihood-based objective, we propose to enhance the algebraic characterization of acyclicity with two adaptive modules for causal structure recovering task: 1) an Asymptotic Causal Mask Learning (ACML) module which leverages learnable priority vectors ($\boldsymbol{p}$) and the Gumbel-Sigmoid function to generate causal masks, ensuring the creation of directed acyclic graphs (DAGs) while optimizing computational efficiency; 2) a Dynamic Graph Parameter Learning (DGPL) module to transform causal learning into decomposed matrix products ($\boldsymbol{W}=\boldsymbol{E}_{s}\boldsymbol{E}^{T}_{t}$), effectively capturing the dynamic causal structure of high-dimensional data and enhancing interpretability. Those leads us can exploit faster gradient-based optimization, such as Adam (Kingma and Ba, 2015), and GPU acceleration.
Contribution. The main contributions of this paper are as follows:
  • To the best of our knowledge, LOCAL is the first attempt to formulate a quasi-maximum likelihood-based score function for learning the dynamic causal structure and shows that more robust and accurate.
  • We proposed two adaptive modules, ACML and DGPL, which further liberate the matrix exponential operations required for causality.
  • We conducted extensive experiments on several synthetic datasets and real-world benchmarks, which proved that LOCAL outperforms state-of-the-art by a significant margin.
2. Related Work
2.1. Score-Based Objective for Causal Discovery
@@
Learning causal structure from data using constraint and score-based techniques is a classic approach (Maxwell Chickering and Heckerman, 1997), but the major drawback is the computational inefficiency. Many recent works have made great efforts to promote computability and practicality: NOTEARS (Zheng et al., 2018) employed a least squares objective to speed up the process, but required augmented Lagrangian to satisfy hard DAG constraints. Based on this, DYNOTEARS (Pamfil et al., 2020) first tried to recover the causal structure of time series data. Golem (Ng et al., 2020) exploits the likelihood-based function to make the objective unconstrained. Those likelihood-based works, however, still struggled with the DAG constraint.
2.2. Constraint Free Causal Discovery
@@
Many efforts have been made to emancipate constraints DAGs, focusing on designing more tractable optimization constraints (Yu et al., 2019; Bello et al., 2022; Cai et al., 2023b; Ferdous et al., 2023) on static data. In contrast to these approaches, Yu et al. (Yu et al., 2021) introduced NOCURL, which separated the topological ordering from the adjacency matrix, thereby eliminating the necessity to optimize the DAG constraint within the objective function. Inspired by NOCURL, a series of causal masks based on topological structures appeared (Ng et al., 2022; Annadani et al., 2023; Massidda et al., 2024). For example, MCSL-MLP (Ng et al., 2022) sampled DAGs from a real matrix $U\in\mathbb{R}^{d\times d}$. BayesDAG (Annadani et al., 2023) inherited NOCRUL’s grad matrix and directly sampled DAGs from the posterior of grad matrix $\boldsymbol{p}(u,v)$. By exploiting the temperature sigmoid function to $\boldsymbol{p}(u,v)$, COSMO (Massidda et al., 2024) provides a fast implementation method when samples DAGs. However, DAG masks for the time domain datasets have rarely been studied.
2.3. Matrix Decomposed for Causal Discovery
@@
The matrix decomposition approach is widely employed to address high-dimensional problems (Fang et al., 2024; Massidda et al., 2024; Chen et al., 2024), with applications spanning large-scale parameter tuning (Hu et al., 2022), image restoration (Chen et al., 2022), and trajectory prediction (Bae et al., 2023). While many of these works focus on learning intrinsic dimensionality to achieve efficient optimization in parameter space, VI-DP-DAG (Charpentier et al., 2022) adopts adjacency matrix decomposition $\boldsymbol{W}$ as $\boldsymbol{W}=\amalg^{T}U\amalg$, while LoRAM (Dong and Sebag, 2023) integrates low-rank matrix factorization with sparsification mechanisms for continuous optimization of DAGs. Those works do not conduct an in-depth exploration of the physical meaning of the decomposition matrix, and the decomposition form is not concise enough.
@@
3. LOCAL: Learning Causal Discovery from Time Series Data
3.1. Problem Definition
@@
We target the causal discovery from a dynamic system. Consider multitudinous stationary time series that contains $d$ causal univariate time series represented as $\mathcal{X}:=[X^{i}_{t}]\in\mathbb{R}^{T\times d}$, where $X^{i}_{t}$ is the recording of $i$-th component of $X_{t}$ at time step $t$, our target is to simultaneously infer the instantaneous dependency $\boldsymbol{W}^{i}$ = $\boldsymbol{W}^{i,j}:X^{i}_{t}\rightarrow X^{j}_{t}|i,j\in 1,\cdots,d;i\neq j\}$ and the lagged dependency $\boldsymbol{A}^{i}$ = $\boldsymbol{A}_{k}^{i,j}:X^{j}_{t-p}\rightarrow X^{i}_{t}|i,j\in 1,\cdots,d;p\in Z^{+}\}$ of $X^{i}_{t}$, from which we may recovery the dynamic causal structure. Following the practice in the dynamic causal structure discovery, we formulate the problem of learning a dynamic Bayesian network (DBN) that captures the dependencies in the $X^{i}_{t}$:
(1) $\begin{gathered}X_{t}=X_{t}\boldsymbol{W}+Y\boldsymbol{A}+N\\ s.t.\text{ }h(\boldsymbol{W})=0\end{gathered}$
@@
3.2. Quasi-Maximum Likelihood-Based Objective
Due to the hard acyclicity constraint (Pamfil et al., 2020), learning a DAG equivalent to the ground truth DAG through directly optimizing Eq. (1) is difficult. To circumvent this issue, we reformulate a score-based method to maximize the data quasi-likelihood of a dynamic linear Gaussian model.
@@ 3.2. Quasi-Maximum Likelihood-Based Objective
(2) $\begin{gathered}\min_{\boldsymbol{W},\boldsymbol{A}}\mathcal{S}(\boldsymbol{W},\boldsymbol{A};X_{t})=\mathcal{L}(\boldsymbol{W},\boldsymbol{A};X_{t})\\ +\lambda_{1}h(\boldsymbol{W})+\lambda_{2}(\mathcal{R}_{sparse}(\boldsymbol{A})+\mathcal{R}_{sparse}(\boldsymbol{W}))\end{gathered}$
@@
Suppose $N$ follows the multivariate normal distribution $\mathcal{N}(0,\sigma_{i}^{2}I)$. Further assumes that the noise variances are equal, i.e. $\sigma_{1}^{2}=\cdots=\sigma_{d}^{2}=\sigma^{2}$. Let $X^{-i}_{t}=(X^{j}_{t},j\neq i)^{T}$ denote the responses of all variables except for the $i$-th variable. Motivated by the least squares estimation (LSE) method (Zhu et al., 2020), we have $\tilde{X}^{i}_{t}=E\{X^{i}_{t}|X^{-i}_{t}\}=\mu_{i}+\sum_{j\neq i}\alpha_{i,j}(X^{j}_{t}-\mu_{j})$, where
(3) $\alpha_{ij}=\frac{(w_{ij}+w_{ji})-\sum_{k}w_{ki}w_{kj}}{1+\sum_{k}w^{2}_{ki}}$
@@
(4) $\tilde{X}^{i}_{t}-\mu_{i}$ $=(1+\|\boldsymbol{W}_{\cdot i}\|^{2})^{-1}(W_{\cdot i}+\boldsymbol{W}^{T}_{\cdot i}-\boldsymbol{W}^{T}_{\cdot i}W_{ii})(x_{t}-\mu)$
@@
By defining $\boldsymbol{D}:=\{\boldsymbol{I}+diag(\boldsymbol{W}^{T}\boldsymbol{W})\}^{-1}$ and $\boldsymbol{S}:=\boldsymbol{I}-\boldsymbol{W}$, we could verify the least squares objective function as
(5) $\frac{1}{n}\sum^{d}_{i=1}\|X^{i}_{t}\boldsymbol{S}-Y^{i}\boldsymbol{A}\|^{2}=\frac{1}{n}\|\{X_{t}\boldsymbol{S}-Y\boldsymbol{A}\}\boldsymbol{D}\boldsymbol{S}^{T}\|^{2}$
@@
(6) $\mathcal{L}(\boldsymbol{W},\boldsymbol{A};X_{t})$ $=\frac{1}{2}\log\sigma^{2}-\log|\text{det}(\boldsymbol{S})|$
$-\frac{1}{2}\frac{(X_{t}\boldsymbol{S}-Y\boldsymbol{A})^{\prime}(X_{t}\boldsymbol{S}-Y\boldsymbol{A})}{\sigma^{2}}$
@@
(7) $\hat{\sigma}^{2}=\frac{1}{n}[X_{t}\boldsymbol{S}-Y\boldsymbol{A}]^{\prime}[X_{t}\boldsymbol{S}-Y\boldsymbol{A}]$
@@
(8) $\mathcal{L}(\boldsymbol{W},\boldsymbol{A};X_{t})=\frac{d}{2}\log(\|X_{t}\boldsymbol{S}-Y\boldsymbol{A}\|^{2})-\log|\text{det}(\boldsymbol{S})|$
Note that, the $\|X_{t}\boldsymbol{S}-Y\boldsymbol{A}\|^{2}$ term in Eq. (8) is the least squares objective, replace it with Eq. (5), we could verify the objective function as
(9) $\begin{gathered}\mathcal{L}(\boldsymbol{W},\boldsymbol{A};X_{t})=\frac{d}{2}\log(\|\{X_{t}\boldsymbol{S}-Y\boldsymbol{A}\}\boldsymbol{D}\boldsymbol{S}^{T}\|^{2})-\log|\text{det}(\boldsymbol{S})|\end{gathered}$
3.3. Asymptotic Causal Mask Learning
Although QMLE can be statistically efficient, the computational cost could be expensive due to the matrix exponential operation of $h(\cdot)$ function (Zheng et al., 2018). Most recent work in causal structure learning deploys an alternative continuous DAG constraint $h(\boldsymbol{W})$ to characterize acyclic graphs. Although decoupling the causal structure into a causal mask and a real matrix may help to give more explanatory results for instantaneous causal effects, it means that the matrix needs to learn changes from $d\times d$ to $d\times(d+1)$. To solve the issue, we propose an Asymptotic Causal Mask Learning (ACML) module to asymptotic infer the causal relationship from data automatically. The ACML module initializes a learnable priority vector $\boldsymbol{p}\in\mathbb{R}^{d}$, which is initialized as a vector with all elements set to 1, and a strictly positive threshold $\omega>0$. Then the smooth orientation matrix of the strict partial order is the binary orientation matrix $\boldsymbol{M}_{\omega}(\boldsymbol{p})_{u,v}\in\{0,1\}^{d\times d}$ such that:
(10) $\begin{gathered}\boldsymbol{M}_{\tau,\omega}(\boldsymbol{p})_{u,v}=\sigma_{\tau,\omega}(\boldsymbol{p}_{v}-\boldsymbol{p}_{u})\end{gathered}$

### v3

@@
Abstract
Discovering the underlying Directed Acyclic Graph (DAG) from time series observational data is highly challenging due to the dynamic nature and complex nonlinear interactions between variables. Existing methods typically search for the optimal DAG by optimizing an objective function but face scalability challenges, as their computational demands grow exponentially with the dimensional expansion of variables. To this end, we propose LOCAL, a highly efficient, easy-to-implement, and constraint-free method for recovering dynamic causal structures. LOCAL is the first attempt to formulate a quasi-maximum likelihood-based score function for learning the dynamic DAG equivalent to the ground truth. Building on this, we introduce two adaptive modules that enhance the algebraic characterization of acyclicity: Asymptotic Causal Mask Learning (ACML) and Dynamic Graph Parameter Learning (DGPL). ACML constructs causal masks using learnable priority vectors and the Gumbel-Sigmoid function, ensuring DAG formation while optimizing computational efficiency. DGPL transforms causal learning into decomposed matrix products, capturing dynamic causal structure in high-dimensional data and improving interpretability. Extensive experiments on synthetic and real-world datasets demonstrate that LOCAL significantly outperforms existing methods and highlight LOCAL’s potential as a robust and efficient method for dynamic causal discovery.
keywords:
@@ Abstract
journal: Applied Soft Computing
[…]
label1label1footnotetext: WeiWei Xinga and Yue Cheng are the Co-Corresponding author. E-mail: wwxing@bjtu.edu.cn, yuecheng@bjtu.edu.cn
@@
Exploration of the underlying causal generation process of dynamic systems is an important task [11, 19] for trustworthy machine learning. Unfortunately, it is unethical, impossible due to technical reasons, or expensive to conduct intervention experiments on the dynamic systems of certain domains [25, 6]. Another challenge is to infer about the structure which may be high dimensional and nonlinear. Some recent works [34, 37, 18, 15] have made significant efforts by employing dynamic Bayesian networks (DBNs) with observational and interventional data: among dynamic systems, as illustrated in Figure 1. The variable $x_{i}$ at timestep $t$ is affected by which variables $x_{j}$ at the same time step (instantaneous dependency) and which variables $x_{j}$ at the previous timestep (lagged dependency)? This question highlights the crucial roles of those algorithms in the interpretable performance of the trained models.
@@
Figure 1: Illustration of instantaneous dependencies (solid lines) and lagged dependencies (dashed lines) in a dynamic Bayesian network (DBN) with $d=3$ nodes and an autoregression order of $p=2$. For clarity, edges that do not influence the variables at time $t$ are shown in a lighter shade.
@@
In order to study the nonparametric DBN, DYNO [34] (i.e., a score-based approach to learning DBNs with a differentiable DAG constraint [43]) was proposed as a proxy to capture the parents of child variables. However, in Section 3.2, our practical analysis shows that the DYNO algorithm and its extensions [37, 18, 15] adopting matrix exponential constraints require an extremely long time to optimize in high-dimensional dynamic systems, even if they smartly adopt interventional data to enhance the identifiability [26, 27]. Then, it is natural to need a proxy model that can infer dynamic causal structures in high-dimensional situations faster, which is also the main goal of our work.
@@
Recently, much of the research on dynamic causal structure learning has concentrated on applying soft sparsity and DAG constraints. For instance, Golem [32] formulated a likelihood-based score function for handling the causality of thousands of nodes. Yu et al. [42] extended it for the recovery of dynamic causal structure. Concurrently, [16] verified the feasibility of further accelerating the learning of DAG based on this likelihood function both theoretically and experimentally. These approaches are aimed at enhancing flexibility and scalability in high dimensions and circumventing rigid structural constraints.
@@
Building on these insights, we propose a novel framework for Learning with Orientation matrix to infer CAusaL structure from time series data, which we call LOCAL. We develop a quasi-maximum likelihood-based dynamic structure learning method with identifiability guarantee. Powered by this quasi-maximum likelihood-based objective, we propose to enhance the algebraic characterization of acyclicity with two adaptive modules for causal structure recovering task: 1) an Asymptotic Causal Mask Learning (ACML) module which leverages learnable priority vectors ($\boldsymbol{p}$) and the Gumbel-Sigmoid function to generate causal masks, ensuring the creation of directed acyclic graphs (DAGs) while optimizing computational efficiency; 2) a Dynamic Graph Parameter Learning (DGPL) module to transform causal learning into decomposed matrix products ($\boldsymbol{W}=\boldsymbol{E}_{s}\boldsymbol{E}^{T}_{t}$), effectively capturing the dynamic causal structure of high-dimensional data and enhancing interpretability. Those leads us to exploit faster gradient-based optimization, such as Adam [24], and GPU acceleration.
Contribution. The main contributions of this work are as follows:
  • 1.
    To the best of our knowledge, this work represents the first principled formulation of a quasi-maximum likelihood score function for dynamic causal structure learning. Our theoretical analysis and empirical validation demonstrate its superior robustness and accuracy compared to existing approaches.
  • 2.
    We propose LOCAL, an innovative framework that integrates two adaptive modules, ACML and DGPL, effectively eliminating the need for matrix exponential operations in causal learning.
  • 3.
    We conduct extensive experiments on both synthetic datasets and real-world benchmarks. On a synthetic dataset with 100 nodes, our method achieves impressive instantaneous and lagged F1 scores of 0.81 and 0.59, respectively.
The remainder of this paper is structured as follows. In Section. 1, we discuss the challenges of dynamic causal discovery and the limitations of existing methods. In Section. 2, we introduce our quasi-maximum likelihood-based objective function and detail the design of the ACML and DGPL modules for learning dynamic DAG structures and improving nonlinear modeling. In Section. 3, we present experimental results demonstrating the advantages of LOCAL in accuracy, efficiency, and interpretability across synthetic, NetSim, and real-world datasets. We also perform ablation studies and parameter analysis. Finally, in Section. 4 and Section. 5, we summarize efficiency of LOCAL and discuss potential future extensions.
1 Related Work
1.1 Score-Based Objective for Causal Discovery
@@
Learning causal structure from data using constraint and score-based techniques was a classic approach [29], but the major drawback was the computational inefficiency. Many recent works have contributed to promoting computability and practicality: NOTEARS [43] employed a least squares objective to speed up the process but required an augmented Lagrangian to satisfy hard DAG constraints. Based on this, DYNO [34] first tried to recover the causal structure of time series data. Golem [32] exploited the likelihood-based function to make the objective unconstrained. Those likelihood-based works, however, still struggled with the DAG constraint.
1.2 Constraint Free Causal Discovery
@@
Many efforts have been made to emancipate constraints DAGs, focusing on designing more tractable optimization constraints [40, 3, 7, 17] on static data. In contrast to these approaches, Yu et al. [41] introduced NOCURL, which separated the topological ordering from the adjacency matrix, thereby eliminating the necessity to optimize the DAG constraint within the objective function. Inspired by NOCURL, a series of causal masks based on topological structures appeared [33, 1, 28]. For example, MCSL-MLP [33] sampled DAGs from a real matrix $U\in\mathbb{R}^{d\times d}$. BayesDAG [1] inherited NOCRUL’s grad matrix and directly sampled DAGs from the posterior of grad matrix $\boldsymbol{p}(u,v)$. By exploiting the temperature sigmoid function to $\boldsymbol{p}(u,v)$, COSMO [28] provided a fast implementation method when sampled DAGs. However, DAG masks for the time domain datasets have rarely been studied.
1.3 Matrix Decomposition for Causal Discovery
@@
The matrix decomposition approach was widely employed to address high-dimensional problems [16, 28, 10], with applications spanning large-scale parameter tuning [22], image restoration [9], and trajectory prediction [2]. While many of these works focused on learning intrinsic dimensionality to achieve efficient optimization in parameter space, VI-DP-DAG [8] adopted adjacency matrix decomposition $\boldsymbol{W}$ as $\boldsymbol{W}=\amalg^{T}U\amalg$, while LoRAM [14] integrated low-rank matrix factorization with sparsification mechanisms for continuous optimization of DAGs. Those works did not conduct an in-depth exploration of the physical meaning of the decomposition matrix, and the decomposition form was not concise enough.
@@
2 LOCAL: Learning Causal Discovery from Time Series Data
2.1 Problem Definition
@@
We target the problem of causal discovery from a dynamic system. Consider multidimensional stationary time series that contains $d$ causal univariate time series represented as $\mathcal{X}:=[X^{i}_{t}]\in\mathbb{R}^{T\times d}$, where $X^{i}_{t}$ is the recording of $i$-th component of $X_{t}$ at time step $t$. Our target is to simultaneously infer the instantaneous dependency $\boldsymbol{W}^{i}$ = $\boldsymbol{W}^{i,j}:X^{i}_{t}\rightarrow X^{j}_{t}|i,j\in 1,\cdots,d;i\neq j\}$ and the lagged dependency $\boldsymbol{A}^{i}$ = $\boldsymbol{A}_{k}^{i,j}:X^{j}_{t-p}\rightarrow X^{i}_{t}|i,j\in 1,\cdots,d;p\in Z^{+}\}$ of $X^{i}_{t}$, from which we may recovery the dynamic causal structure. Following the practice in the dynamic causal structure discovery, we formulate the problem of learning a dynamic Bayesian network (DBN) that captures the dependencies in the $X^{i}_{t}$:
$\begin{gathered}X_{t}=X_{t}\boldsymbol{W}+Y\boldsymbol{A}+N\\ s.t.\text{ }h(\boldsymbol{W})=0\end{gathered}$ (1)
@@
2.2 Quasi-Maximum Likelihood-Based Objective
Due to the hard acyclicity constraint [34], learning a DAG equivalent to the ground truth DAG through directly optimizing Eq. (1) is difficult. To circumvent this issue, we reformulate a score-based method to maximize the data quasi-likelihood of a dynamic linear Gaussian model.
@@ 2.2 Quasi-Maximum Likelihood-Based Objective
$\begin{gathered}\min_{\boldsymbol{W},\boldsymbol{A}}\mathcal{S}(\boldsymbol{W},\boldsymbol{A};X_{t})=\mathcal{L}(\boldsymbol{W},\boldsymbol{A};X_{t})\\ +\lambda_{1}h(\boldsymbol{W})+\lambda_{2}(\mathcal{R}_{sparse}(\boldsymbol{A})+\mathcal{R}_{sparse}(\boldsymbol{W}))\end{gathered}$ (2)
@@
Suppose $N$ follows the multivariate normal distribution $\mathcal{N}(0,\sigma_{i}^{2}I),\,i=1,2,\dots,d$. Further assumes that the noise variances are equal, i.e. $\sigma_{1}^{2}=\cdots=\sigma_{d}^{2}=\sigma^{2}$. Let $X^{-i}_{t}=(X^{j}_{t},j\neq i)^{T}$ denote the responses of all variables except for the $i$-th variable. Motivated by the least squares estimation (LSE) method [45], we have $\tilde{X}^{i}_{t}=E\{X^{i}_{t}|X^{-i}_{t}\}=\mu_{i}+\sum_{j\neq i}\alpha_{i,j}(X^{j}_{t}-\mu_{j})$, where
$\alpha_{ij}=\frac{(w_{ij}+w_{ji})-\sum_{k}w_{ki}w_{kj}}{1+\sum_{k}w^{2}_{ki}}$ (3)
@@
$\tilde{X}^{i}_{t}-\mu_{i}$ $=(1+\|\boldsymbol{W}_{\cdot i}\|^{2})^{-1}(W_{\cdot i}+\boldsymbol{W}^{T}_{\cdot i}-\boldsymbol{W}^{T}_{\cdot i}W_{ii})(x_{t}-\mu)$ (4)
@@
By defining $\boldsymbol{D}:=\{\boldsymbol{I}+diag(\boldsymbol{W}^{T}\boldsymbol{W})\}^{-1}$ and $\boldsymbol{S}:=\boldsymbol{I}-\boldsymbol{W}$, we can verify the least squares objective function as
$\frac{1}{n}\sum^{d}_{i=1}\|X^{i}_{t}\boldsymbol{S}-Y^{i}\boldsymbol{A}\|^{2}=\frac{1}{n}\|\{X_{t}\boldsymbol{S}-Y\boldsymbol{A}\}\boldsymbol{D}\boldsymbol{S}^{T}\|^{2}$ (5)
@@
$\mathcal{L}(\boldsymbol{W},\boldsymbol{A};X_{t})$ $=\frac{1}{2}\log\sigma^{2}-\log|\text{det}(\boldsymbol{S})|-\frac{1}{2}\frac{(X_{t}\boldsymbol{S}-Y\boldsymbol{A})^{\prime}(X_{t}\boldsymbol{S}-Y\boldsymbol{A})}{\sigma^{2}}$ (6)
@@
$\hat{\sigma}^{2}=\frac{1}{n}[X_{t}\boldsymbol{S}-Y\boldsymbol{A}]^{\prime}[X_{t}\boldsymbol{S}-Y\boldsymbol{A}]$ (7)
@@
$\mathcal{L}(\boldsymbol{W},\boldsymbol{A};X_{t})=\frac{d}{2}\log(\|X_{t}\boldsymbol{S}-Y\boldsymbol{A}\|^{2})-\log|\text{det}(\boldsymbol{S})|$ (8)
Note that the $\|X_{t}\boldsymbol{S}-Y\boldsymbol{A}\|^{2}$ term in Eq. (8) is the least squares objective, replace it with Eq. (5), and we could verify the objective function as
$\begin{gathered}\mathcal{L}(\boldsymbol{W},\boldsymbol{A};X_{t})=\frac{d}{2}\log(\|\{X_{t}\boldsymbol{S}-Y\boldsymbol{A}\}\boldsymbol{D}\boldsymbol{S}^{T}\|^{2})-\log|\text{det}(\boldsymbol{S})|\end{gathered}$ (9)
2.3 Asymptotic Causal Mask Learning
Although QMLE can be statistically efficient, the computational cost could be expensive due to the matrix exponential operation of $h(\cdot)$ function [43]. Most recent work in causal structure learning deploys an alternative continuous DAG constraint $h(\boldsymbol{W})$ to characterize acyclic graphs. Decoupling the causal structure into a causal mask and a real matrix can improve explainability for instantaneous causal effects. However, this requires the matrix to adapt from $d\times d$ to $d\times(d+1)$. To address this issue, we propose the Asymptotic Causal Mask Learning (ACML) module , which asymptotically infers causal relationships from data. The ACML module initializes a learnable priority vector $\boldsymbol{p}\in\mathbb{R}^{d}$, which is initialized as a vector with all elements set to 1, and a strictly positive threshold $\omega>0$. The smooth orientation matrix of the strict partial order is given by the binary orientation matrix $\boldsymbol{M}_{\omega}(\boldsymbol{p})_{u,v}\in\{0,1\}^{d\times d}$ such that:
$\begin{gathered}\boldsymbol{M}_{\tau,\omega}(\boldsymbol{p})_{u,v}=\sigma_{\tau,\omega}(\boldsymbol{p}_{v}-\boldsymbol{p}_{u})\end{gathered}$ (10)